Mr Garcia had 2/3 pumpkin pie left. His sons ate 3/8 of what was left. What fraction of the whole pie did his sons eat?
Answer by stanbon(75887)

Mr Garcia had 2/3 pumpkin pie left. His sons ate 3/8 of what was left. What fraction of the whole pie did his sons eat?
Ans: (3/8)(2/3) = 2/8 = 1/4
------------
Cheers,
Stan H.
-----------